OpenStudy (anonymous):

Solve 4(c - 3) = 4c + 8.

OpenStudy (anonymous):

4(c - 3) = 4c + 8 4c - 12 = 4c + 8 +12 +12 4c = 4c + 20 -4c -4c 0 = 20 No solution